OpenStudy (anonymous):

Factor completely: 4x3 + 28x2 + 7x + 49 (x − 7)(4x2 + 7) (x − 7)(4x2 − 7) (x + 7)(4x2 + 7) (x + 7)(4x2 − 7)

jimthompson5910 (jim_thompson5910):

4x^3 + 28x^2 + 7x + 49 4x^2(x+7) + 7(x+7) (4x^2+7)(x+7) (x+7)(4x^2+7)

OpenStudy (anonymous):

4x^2(x+7)+7(x+7) (4x^2+7)(x+7)

OpenStudy (anonymous):

Simplifying 12(x2 + 3x + -18) = 0 Reorder the terms: 12(-18 + 3x + x2) = 0 (-18 * 12 + 3x * 12 + x2 * 12) = 0 (-216 + 36x + 12x2) = 0 Solving -216 + 36x + 12x2 = 0 Solving for variable 'x'. Factor out the Greatest Common Factor (GCF), '12'. 12(-18 + 3x + x2) = 0 Factor a trinomial. 12((-6 + -1x)(3 + -1x)) = 0 Ignore the factor 12. Subproblem 1Set the factor '(-6 + -1x)' equal to zero and attempt to solve: Simplifying -6 + -1x = 0 Solving -6 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '6' to each side of the equation. -6 + 6 + -1x = 0 + 6 Combine like terms: -6 + 6 = 0 0 + -1x = 0 + 6 -1x = 0 + 6 Combine like terms: 0 + 6 = 6 -1x = 6 Divide each side by '-1'. x = -6 Simplifying x = -6 Subproblem 2Set the factor '(3 + -1x)' equal to zero and attempt to solve: Simplifying 3 + -1x = 0 Solving 3 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-3' to each side of the equation. 3 + -3 + -1x = 0 + -3 Combine like terms: 3 + -3 = 0 0 + -1x = 0 + -3 -1x = 0 + -3 Combine like terms: 0 + -3 = -3 -1x = -3 Divide each side by '-1'. x = 3 Simplifying x = 3Solutionx = {-6, 3}

OpenStudy (anonymous):

bloog it never said it was an equation. just something to factor. there is nothing to solve for.